If you lift a box using a 100 J of work within 5 s, how much power has been done? How much is the power if it’s lifted using a 10 J of work?



1
Expert's answer
2021-12-01T10:28:45-0500

(a)

"P=\dfrac{W}{t}=\dfrac{100\ J}{5\ s}=20\ W."

(b)

"P=\dfrac{W}{t}=\dfrac{10\ J}{5\ s}=2\ W."